OpenStudy (anonymous):

If you roll a number cube 12 times, about how many times would you expect to roll a 5 or 6? A. 2 B. 5 C. 4 D. 10

OpenStudy (anonymous):

merry, u hav any guess on how to do this one?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

a dice hav 6possible outcome. so probability 5or6 occur: p(5 or 6)= 1/6+1/6= 1/3 since it rolled 12times, expected number of 5or6 occur is: E(5 or 6)= 1/3*12= 4

OpenStudy (anonymous):

thats exactly what i guessed